Susan Wilson is saving to buy a house in five years. She plans to put 20 percent down at that time, and she believes that she will need $27,000 for the down payment. If Susan can invest in a fund that pays 6.40 percent annual interest, compounded quarterly, how much will she have to invest today to have enough money for the down payment?
Answer #1

Here we will use the following formula:

FV = PV * (1 + r%)n

where, FV = Future value = $27000, PV = Present value, r = rate of interest = 6.4% compounded quarterly. So, Quarterly rate = 6.4% / 4 = 1.6%, n= time period = 5 * 4 = 20 Quarter years.

now, putting theses values in the above equation, we get,

$27000 = P * (1 + 1.6%)20

$27000 = P * (1 + 0.016)20

$27000 = P * (1.016)20

$27000 = P * 1.37364389069

P = $27000 / 1.37364389069

P = $19655.75

So, she will have to invest $19655.75 today.
Answer #2

To calculate the amount Susan needs to invest today to have enough money for the down payment, we can use the formula for compound interest:

A = P(1 + r/n)^(nt)

Where: A = Amount accumulated after t years P = Principal amount (initial investment) r = Annual interest rate (as a decimal) n = Number of compounding periods per year t = Number of years

In this case, Susan wants to have $27,000 for the down payment in five years, and she can invest in a fund that pays 6.40% annual interest, compounded quarterly.

Let's calculate the amount she needs to invest (P):

A = $27,000 (down payment amount) r = 6.40% = 0.064 (annual interest rate) n = 4 (quarterly compounding) t = 5 (number of years)

Using the formula, we can rearrange it to solve for P:

P = A / (1 + r/n)^(nt)

P = $27,000 / (1 + 0.064/4)^(4*5)

P ≈ $20,095.25

Therefore, Susan needs to invest approximately $20,095.25 today to have enough money for the down payment of $27,000 in five years, assuming a 6.40% annual interest rate compounded quarterly.


answered by: mervetokaz